# Adaptive Gait Modeling and Optimization for Principally Kinematic Systems

### v1

@@ I Introduction
The field of robotics has experienced significant advancements over the past few decades, evolving from systems designed for explicit situations into more generalizable and versatile machines. This transformation has been driven by the ever-growing demands of robots that are capable of interacting with dynamic and unstructured environments. In many such environments, learning from in-situ experiences is crucial to functioning in the field. These challenges are particularly evident for systems with high internal degrees of freedom and complex system-environment interactions, where subtle changes in environmental characteristics can drive fundamental differences in emergent behavior.
@@
Current paradigms of motion control are not ideally suited to attaining precision in these unstructured settings. Traditional bottom-up, model predictive approaches to robotic control heavily rely on explicit accounting of physical quantities which can be derived from first principles or hand-crafted to the desired precision of the user [1, 2]. Adaptive system identification extends models to react to anticipated changes in the model. The rate and quality of adaptation are affected strongly by the assumptions about observable perturbations of the model [3, 4, 5]. In general, these methods are ineffective when either (a) the range of assumed adaptations is so large that overfitting occurs or (b) the assumptions are too specific such that violations of the assumptions are common. More recent approaches to motion control include reinforcement learning and deep learning approaches, both of which require large magnitudes of training data or infrastructurally cumbersome quantities of real-world data [6, 7, 8]. Here, the robot is particularly ill-suited to succeed in unanticipated scenarios, since (by definition) they would not be included in the training set. Efforts in domain randomization have yielded reinforcement learning agents more capable of generalization and sim-to-real transfer, but often at the expense of achieving precise mastery of individual domains [9, 10]. Learning frameworks such dynamic mixtures of experts [11, 12] attempt to achieve robustness and precision with distributed expertise, but the computation for these experts to both decouple from each others’ domains and refine individual expertise has not yet been designed for real-time adaptation in complex fielded environments. Adaptive control, reinforcement learning, and deep learning architectures are not designed to rapidly and precisely handle unanticipated changes to the dynamics.
@@
This limits our ability to advance robots to new application spaces that can handle complexities such as customized body parts, imprecise fabrications, and poorly characterized environments. Such considerations are crucial for soft robotics [13, 14], nano-robotics [15], medical robots [16], and bio-hybrid robots [17], as each is subject to these considerations. In this work, we use an adaptive system identification framework to leverage the top-down modeling insights of geometric mechanics [18]. This permits a framework for adaptation that captures the general space of possible dynamics changes within a class of physics, but it is representationally compact enough to be fit rapidly to data in situ. The result enables adapting motion models to distinctly new regimes without knowing a priori the nature of that change (e.g. a substrate change, morphology change, or weight distribution change). The assumptions made are that the dissipative nature of the physics persists and that the recent history of the behavior is evolving in a homogeneous environment. A key contribution of our work is that empirical modeling is moved online as the system operates and captures a general class of potentially time-varying factors such as environments and payload changes, affording effective motion control despite unanticipated changes to the robot morphology, weight distribution, or substrate conditions.
@@
We restrict our attention in this work to systems that are principally kinematic [19], a property developed in the geometric mechanics community that describes the locomotion of systems governed by Rayleigh dissipation or conservation of zero momentum. These systems enjoy an exploitable reduction from second-order mechanical dynamics to ones that possess a first-order structure, namely a linear mapping from internal configuration velocity to body velocity. Recent work has shown that this reduction provides a reduced computational structure for fitting robot models that are broadly descriptive of many classes of animal and robot behaviors [20, 21]. These works have established that many robot and animal behaviors are describable by the space of Rayleigh dissipation models, which span a dense and broad variety of animal morphologies, weight distributions, and environmental substrates. Bittner et al. [18] introduced a data-driven method for constructing a local model in the neighborhood of an observed gait. This method relies on data points collected from stochastically perturbed and iteratively repeated behaviors. The ability to construct local models makes it possible to sample candidate gaits offline, which provides an opportunity for sample efficient gait optimization as well as gait modulations for continuous gait control [22].
@@
Here, we capitalize on the use of adaptive filters in constructing the aforementioned local models in real-time. The advantages of this reframing are multifold. In this continuously updating paradigm, the model automatically adapts to changes in physical interactions dynamically, rather than awaiting a batch update as in prior work. This actively updated model enables tracking of the model confidence, which is crucial in determining how much empirical data is required to obtain a model with acceptable prediction quality for behavior refinement and control. We show that the rate and quality of adaptation permit refinements of behaviors to master unanticipated environments with speed and precision. We discuss that as internal complexity is added to the system, we encounter situations where engineering decisions between speed and reliable precision need to be made.
@@ II-A Principally Kinematic Systems
In this work, we will pursue an adaptation scheme that applies to a subclass of these systems. Here, we focus on principally kinematic systems. Commonly, these can be observed as systems where Rayleigh dissipation dominates the dynamics [26]. Recent work has suggested this class of dissipation model, despite its simplicity, is highly descriptive of a broad class of robots and animals [20, 21]. Principally kinematic systems are also observed where conservation of zero momentum drives the physics [27]. These scenarios are less commonly observed but applicable to orbital manipulation. In principally kinematic systems, the equations of motion can be reduced, resulting in a shape-defined linear mapping from shape velocity to body velocity. The resulting model is
@@ II-B Data-driven methods for geometric systems
The initial limitations of this approach involved its applicability to systems with momentum. Follow up work was done to capture and model the behavior of systems with dissipating momentum while maintaining a first-order prediction structure [28]. Additional follow-on work extended these models to principally kinematic systems that do not possess high bandwidth control of their shape dynamics. Initial extensions addressed systems with high bandwidth control of a subset of the entire shape space [29] but eventually worked towards addressing systems with low bandwidth control of shape variables [30]. In this paper, we explore an extension to the initial work [18] that permits real-time updates for in-situ modeling and adaptation in the field. Simultaneously, we work directly towards the overall architecture that can support these extensions into adjacent geometric domains.
@@ III-A Adaptive Filters
Our method employs adaptive filters, which are widely used in system identification. These filters are designed to estimate the parameters of a linear model by recursively updating the weights as new data samples are available. The main usage of adaptive filters in system identification is that they can track time-varying systems and adapt to changes in the underlying process. Recursive Least Square (RLS) filters are a class of adaptive filters that specifically minimize the mean square error of the predicted filter output and the actual system output (measurements) [31]. Instead of collecting samples in batches and then fitting a model, we designed a cluster of adaptive filters on top of the data-driven modeling framework so that the model can be updated simultaneously as samples are collected. RLS filters fit the role directly in the linear regression problem in Eq. 3.
@@ III-B Model Metrics
Here, is the forgetting factor, which is set to be nearly 1. This number determines how heavily new samples are weighted towards the measure and how fast the history of the metric fades away. Similar to the forgetting factor in the RLS algorithm, (together with other parameters) effectively controls how much history the metric is carrying. For convenience, we will later refer a relatively ’rapid’ forgetting factor to ’s further away from 1, and a ’slow’ forgetting factor to ’s closer to 1. In practice, we try to match this metric memory to the effective model memory, in terms of the number of cycles of empirical data, so that it reflects model qualities on the same amount of data. Note that, in doing so, the two forgetting factors are going to be different because each RLS filter updates only when the sample phase falls into its window while the metric is being updated with every sample collected.
@@ IV Accuracy of Adaptive Geometric Modeling
We first seek to demonstrate the speed and precision through which RLS can provide geometric models of locomotion in real-time. We demonstrate the method’s performance first on a toy locomotion system, the Purcell swimmer [32], whose local connection can be analytically derived [33, 34]. This analytical system provides a test bed for our method that has clear foundations and has been well explored in the robotics literature. As noted earlier, small soft swimming machines are included within the principally kinematic operating environment for which these methods are designed.
We compared our adaptive models to the batch model mentioned in II-B, in terms of model metrics covered in III-B. Again, captures the knowledge of the model beyond (the phase-averaged behavior) up to (full reconstruction of the ground truth). We generated trials of cycles of noisy shape trajectories as mentioned in III-C. We split the data in half, selecting two cycle trials at a time (100 pairs) and evaluating the model computed on the training set against the held out cycle test set. In Fig. 1, the model accuracies are plotted with respect to the robot’s experience throughout the trial. This informs us how the prediction quality grows as the robot’s experience grows. The model metrics are shown in boxplots where each box corresponds to percentiles of a model trained with the first cycles of data collected and tested on the holdout set. When learning from no history, the adaptive model is capable of obtaining an accuracy of in 10 cycles, meaning that it can understand forty percent of all perturbations from the limit cycle. This nears the approximate convergent prediction quality of both methods after 40 cycles. The adaptive model metrics are computed continuously, and so are available in between the discretely updated batch model.
The adaptive method averages a higher overall accuracy than the batch method. The variation in model quality over the duration of the 40 cycles for the adaptive method speaks to the overfitting produced by the recency bias of the measurements. Forgetting factors are designed to help weight current (and possibly different) domains more highly, but will also lead the model to potentially overweight data that it has seen more recently. The batch model considers all samples equally, and it is thus less prone to overfitting. Next, we revisit the forgetting factor as a key enabler for the swimmer to regain predictive power on new, unanticipated substrates Fig. 2.
@@ V Adaptation to Changing Environments
In the derivation of the Purcell swimmer’s connection matrix (which defines its mapping from shape velocity to body velocity), the ratio of lateral and longitudinal drag coefficients [35] directly affects the nature of locomotion. Typically, this ratio is assumed to be larger than one because the swimmer links are assumed to be slender such that drag is stronger in the lateral direction of link motion. When the drag ratio is , the system is unable to exploit the asymmetry of drag to create motion and is unable to displace its center of mass. When the drag ratio approaches infinite, each link acts like a non-slip wheel, allowing it to move farther per cycle than when the drag ratio is classically set to .
@@ VII Discussion and Conclusion
In this paper, we illustrate the real-time capabilities of our adaptive methods for constructing local models for in-situ motion control and behavior refinement. One immediate advantage of integrating adaptive features is the model’s capacity to autonomously adjust in real-time to unanticipated events such as variations in terrain or payload. The prior batch modeling was unreliable at achieving this real-time capability due to the lengthy requirement to run a batch udpate at every iteration. This real-time enhancement brings fast, precise in-situ modeling to the field, which has classically been a challenge for platforms with poor first principles models as well as systems that encounter unanticipated scenarios (i.e., not accounted for in training sets). Unlike reinforcement learning agents and finely calibrated, explicit physics models, our adaptive scheme is designed to refine models quickly and accurately with no assumptions about morphology, weight distribution, or environmental characteristics (so long as the behavior remains principally kinematic).
@@
We demonstrated this by showing the ability of the adaptive RLS technique to recapture an accurate physical model under significant substrate variation. The adaptation outperformed prior work in terms of speed and precision on the three-link Purcell swimmer as shown in Figures 1 and 2. These results indicate that the proposed method is an asset for attaining effective motion control in dynamic environments. We noted the relationship between forgetting factors, the rate of adaptation, and the consequences of overfitting as observed in our results. Expectedly, rapid forgetting factors yielded the fastest response to new substrates but the least stable performance. Having a dial to tune can be crucial in settings where either (a) speed is critical to success and risk is tolerable (b) speed is not important and safety is paramount.
@@
We presented an optimization architecture where the system executes stochastically perturbed repetitions of the same gait until it gains requisite confidence in its data-driven model of the local dynamics. When this confidence is met, it leads to high-frequency steps along a computed policy gradient, using this data-driven model to select new gaits that locally improve some performance criteria. This led to a significant performance boost for the state of the art, where the Purcell swimmer optimized its gait to an improvement within 60 cycles in its top quartile of performance. It is likely that a large range of the swimmer behavior is attributed to a high forgetting factor and admissible model prediction threshold , which could interact to give false confidence, resulting in suboptimal steps along the policy gradient. Principled selection of these parameters will likely yield different outcomes, such as a less rapid optimizer that converges with near certainty to a similar behavior. This will be the subject of further research.

### v2

@@ I Introduction
The field of robotics has experienced significant advancements over the past few decades, evolving from systems designed for explicit situations into more generalizable and versatile machines. This transformation has been driven by the ever-growing demand for robots that are capable of interacting with dynamic and unstructured environments. In many such environments, learning from in-situ experiences is crucial to functioning in the field. These challenges are particularly evident for systems with high internal degrees of freedom and complex system-environment interactions, where subtle changes in environmental characteristics can drive fundamental differences in emergent behavior.
@@
Current paradigms of motion control are not ideally suited to attaining precision in these unstructured settings. Traditional bottom-up, model predictive approaches to robotic control heavily rely on explicit accounting of physical quantities which can be derived from first principles or hand-crafted to the desired precision of the user [1, 2]. Adaptive system identification extends models to react to anticipated changes in the model. The rate and quality of adaptation are affected strongly by the assumptions about observable perturbations of the model [3, 4, 5]. In general, these methods are ineffective when either (a) the range of assumed adaptations is so large that overfitting occurs or (b) the assumptions are too specific such that violations of the assumptions are common. More recent approaches to motion control include reinforcement learning and deep learning approaches, both of which require large, infrastructurally cumbersome quantities of training data (simulated or real) [6, 7, 8]. Here, the robot is particularly ill-suited to succeed in unanticipated scenarios, since (by definition) they would not be included in the training set. Efforts in domain randomization have yielded reinforcement learning agents more capable of generalization and sim-to-real transfer, but often at the expense of achieving precise mastery of individual domains [9, 10]. Learning frameworks such dynamic mixtures of experts [11, 12] attempt to achieve robustness and precision with distributed expertise, but the computation for these experts to both decouple from each others’ domains and refine individual expertise has not yet been designed for real-time adaptation in complex fielded environments. Adaptive control, reinforcement learning, and deep learning architectures are not designed to rapidly and precisely handle unanticipated changes to the dynamics.
@@
This limits our ability to advance robots to new application spaces that can handle complexities such as customized body parts, imprecise fabrications, and poorly characterized environments. Such considerations are crucial for soft robots [13, 14], nano-robots [15], medical robots [16], and bio-hybrid robots [17], as each are subject to these considerations. In this work, we use an adaptive system identification framework to leverage the top-down modeling insights of geometric mechanics [18]. This permits a framework for adaptation that captures the general space of possible dynamics changes within a class of physics, but it is representationally compact enough to be fit rapidly to data in situ. The result enables adapting motion models to distinctly new regimes without knowing a priori the nature of that change (e.g. a substrate change, morphology change, or weight distribution change). The assumptions made are that the dissipative nature of the physics persists and that the recent history of the behavior is evolving in a homogeneous environment. The key contribution of our work is that empirical modeling is enhanced to update dynamically in subsecond time with new observations (shape and body velocity), permitting rapid adaptation to abrupt, unanticipated changes in robot morphology, weight distribution, or substrate conditions.
@@
We restrict our attention in this work to systems that are principally kinematic [19], a property developed in the geometric mechanics community that describes the locomotion of systems governed by dissipative forces or conservation of zero momentum. These systems enjoy an exploitable reduction from second-order mechanical dynamics to ones that possess a first-order structure, namely a linear mapping from internal configuration (shape) velocity to body velocity. Recent work has shown that this reduction provides a reduced computational structure for fitting robot models that are broadly descriptive of many classes of animal and robot behaviors [20, 21]. These works have established that many robot and animal behaviors are describable by the space of dissipative models, which span a dense and broad variety of animal morphologies, weight distributions, and environmental substrates. Bittner et al. [18] introduced a data-driven method for constructing a local model in the neighborhood of an observed gait. This method relies on data points collected from stochastically perturbed and iteratively repeated behaviors. The ability to construct local models makes it possible to sample candidate gaits offline, which provides an opportunity for sample efficient gait optimization as well as gait modulations for continuous gait control [22].
@@
Here, we capitalize on the use of adaptive filters in constructing the aforementioned local models in real time. The advantages of this reframing are multifold. In this continuously updating paradigm, the model automatically adapts to changes in physical interactions dynamically, rather than awaiting a batch update as in prior work. This actively updated model enables tracking of the model confidence, which is crucial in determining how much empirical data is required to obtain a model with acceptable prediction quality for behavior refinement and control. We show that the rate and quality of adaptation permit refinements of behaviors to master unanticipated environments with speed and precision. We discuss that as internal complexity is added to the system, we encounter situations where engineering decisions between speed and reliable precision need to be made.
@@ II-A Principally Kinematic Systems
In this work, we will pursue an adaptation scheme that applies to a subclass of these systems. Here, we focus on principally kinematic systems. Commonly, these can be observed as systems where dissipative forces dominate the dynamics [26]. Recent work has suggested this class of dissipation model, despite its simplicity, is highly descriptive of a broad class of robots and animals [20, 21]. Principally kinematic systems are also observed where conservation of zero momentum drives the physics [27]. These scenarios are less commonly observed but applicable to orbital manipulation. In principally kinematic systems, the equations of motion can be reduced, resulting in a shape-defined linear mapping from shape velocity to body velocity. The resulting model is
@@ II-B Data-driven methods for geometric systems
The initial limitations of this approach involved its applicability to systems with momentum. Follow-up work was done to capture and model the behavior of systems with dissipating momentum while maintaining a first-order prediction structure [28]. Additional follow-on work extended these models to principally kinematic systems that do not possess high bandwidth control of their shape dynamics. Initial extensions addressed systems with high bandwidth control of a subset of the entire shape space [29] but eventually worked towards addressing systems with low bandwidth control of shape variables [30]. In this paper, we explore an extension to the initial work [18] that permits real-time updates for in-situ modeling and adaptation in the field. Simultaneously, we work directly towards the overall architecture that can support these extensions into adjacent geometric domains.
@@ III-A Adaptive Filters
Our method employs adaptive filters, which are widely used in system identification. These filters are designed to estimate the parameters of a linear model by recursively updating the weights as new data samples are available. The main usage of adaptive filters in system identification is that they can track time-varying systems and adapt to changes in the underlying process. Recursive Least Square (RLS) filters are a class of adaptive filters that specifically minimize the mean square error of the predicted filter output and the actual system output (measurements) [31]. Instead of collecting samples in batches and then fitting a model, we designed a cluster of adaptive filters on top of the data-driven modeling framework so that the model can be updated simultaneously as samples are collected. RLS filters fit the role directly in the linear regression problem in (3).
@@ III-B Model Metrics
Here, is the forgetting factor, which is set to be nearly 1. This number determines how heavily new samples are weighted towards the measure and how fast the history of the metric fades away. Similar to the forgetting factor in the RLS algorithm, (together with other parameters) effectively controls how much history the metric is carrying. For convenience, we will later refer a relatively ‘rapid’ forgetting factor to ’s further away from 1, and a ‘slow’ forgetting factor to ’s closer to 1.
@@ IV Accuracy of Adaptive Geometric Modeling
We first seek to demonstrate the speed and precision through which RLS can estimate geometric models of locomotion in real time, provided access to ground truth history of shape and body velocity data. We demonstrate the method’s performance first on a toy locomotion system, the Purcell swimmer [32], whose local connection can be analytically derived [33, 34]. This low Reynolds swimmer is a classical example of a principally kinematic system, so its locomotion is explained entirely by our modeling assumptions.
Selecting the right data size for the batch model is important in balancing data efficiency and model accuracy, and it is not easily generalizable for arbitrary systems. One advantage of the adaptive model is that it can keep track of the model’s accuracy in real time so that there is no need to guess a data size as in the batch model. We compared our adaptive models to the batch model mentioned in II-B, in terms of model metrics covered in III-B. Again, captures the knowledge of the model beyond (the phase-averaged behavior) up to (full reconstruction of the ground truth). We generated trials of cycles of noisy shape trajectories as mentioned in III-C. We split the data in half, selecting two cycle trials at a time (100 pairs) and evaluating the model computed on the training set against the held out cycle test set. In Fig. 1, the model accuracies are plotted with respect to the robot’s experience throughout the trial. This informs us how the prediction quality grows as the robot’s experience grows. The model metrics are shown in boxplots where each box corresponds to percentiles of a model trained with the first cycles of data collected and tested on the holdout set. When learning from no history, the adaptive model is capable of obtaining an accuracy of in 5 cycles, meaning that it can understand forty percent of all perturbations from the limit cycle. This nears the approximate convergent prediction quality of both methods after 40 cycles. The adaptive model metrics are computed continuously, and so are available in between the discretely updated batch model.
The adaptive model performs relatively well even when there are relatively few samples, whereas the batch model performs best with large datasets. The variation in model quality over the duration of the 40 cycles for the adaptive method speaks to the overfitting produced by the recency bias of the measurements.
Forgetting factors are designed to help weight current (and possibly different) domains more highly, but will also lead the model to potentially overweight data that it has seen more recently. The batch model considers all samples equally, and it is thus less prone to overfitting. Next, we revisit the forgetting factor as a key enabler for the swimmer to regain predictive power on new, unanticipated substrates Fig. 2.
@@ V Adaptation to Changing Environments
In the derivation of the Purcell swimmer’s connection matrix (which defines its mapping from shape velocity to body velocity), the ratio of lateral and longitudinal drag coefficients [35] directly affects the nature of locomotion. When the drag ratio is , the system is unable to exploit the asymmetry of drag to create motion and is unable to displace its center of mass. When the drag ratio approaches infinite, each link acts like a non-slip wheel, allowing it to move farther per cycle than when the drag ratio is classically set to .
@@ VII Discussion and Conclusion
In this paper, we illustrate the real-time capabilities of our adaptive methods for constructing local models for in-situ motion control and behavior refinement. One immediate advantage of integrating adaptive features is the model’s capacity to autonomously adjust in real-time to unanticipated events such as variations in terrain or payload. The prior batch modeling was unreliable at achieving this real-time capability due to the lengthy requirement to run a batch udpate at every iteration. This real-time enhancement brings fast, precise in-situ modeling to the field, which has classically been a challenge for platforms with poor first principles models as well as systems that encounter unanticipated scenarios (i.e., not accounted for in training sets). Unlike reinforcement learning agents and finely calibrated, explicit physics models, our adaptive scheme is designed to refine models quickly and accurately with no assumptions about morphology, weight distribution, or environmental characteristics (so long as the behavior remains principally kinematic). Without adaptive control, recent work shows that redundancy in the robot body can improve robustness to rugged terrain [36]. A compelling study might involve synthesis of these design principles with the adaptation schemes discussed in this work.
@@
We demonstrated this by showing the ability of the adaptive RLS technique to recapture an accurate physical model under significant substrate variation. The adaptation outperformed prior work in terms of speed and data efficiency on the Purcell swimmer as shown in Figures 1 and 2. These results indicate that the proposed method is an asset for attaining effective motion control in dynamic environments. We noted the relationship between forgetting factors, the rate of adaptation, and the consequences of overfitting as observed in our results. Expectedly, rapid forgetting factors yielded the fastest response to new substrates but the least stable performance. Having a dial to tune can be crucial in settings where either (a) speed is critical to success and risk is tolerable (b) speed is not important and safety is paramount.
@@
We presented an optimization architecture where the system executes stochastically perturbed repetitions of the same gait until it gains requisite confidence in its data-driven model of the local dynamics. When this confidence is met, it leads to high-frequency steps along a computed policy gradient, using this data-driven model to select new gaits that locally improve some performance criteria. This led to a significant performance boost for the state of the art, where the Purcell swimmer optimized its gait to an improvement within 60 cycles in its top quartile of performance. It is likely that a large range of the swimmer behavior is attributed to a high forgetting factor and admissible model prediction threshold , which could interact to give false confidence, resulting in suboptimal steps along the policy gradient. Principled selection of these parameters will likely yield different outcomes, such as a less rapid optimizer that converges with near certainty to a similar behavior.